# Robust Control Lyapunov-Value Functions for Nonlinear Disturbed Systems

### v2

@@ III-A TV-R-CLVF
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}J_{\gamma}(t,x,u(\cdot),\lambda),$ (10)
@@
$J_{\gamma}(t,x,u(\cdot),d(\cdot)))=\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\bigl{(}\xi(s;t,x,u(\cdot),d(\cdot))\bigl{)},$ (11)
@@
The TV-R-CLVF is Lipschitz in $x,t$ and bounded for any open set $\mathcal{C}$
@@
See the Appendix of [33]. ∎
We now present that the TV-R-CLVF satisfies the dynamic programming principle, and is the unique viscosity solution to the TV-R-CLVF VI.
@@
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V_{\gamma}(\xi(t+\delta),t+\delta),$
@@
The proofs of the above two Theorems can be obtained analogously following Theorems 2 and 3 in [29], and are omitted here. It should be noted that the definitions of viscosity solutions for the terminal value problem (TVP) and initial value problem (IVP) are slightly different. For more details, the reader is referred to [15, 16] for IPV and [34] for TVP.
With Theorems 1 and 2, the TV-R-CLVF can be again obtained by solving equation (13) for $s\in[t,0]$. Note that in equation (13), define $H:\mathcal{D}_{\gamma}\times\mathbb{R}^{n}\mapsto\mathbb{R}$,
$H(x,p)=\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}p\cdot f(x,u,d).$
$H$ is called the Hamiltonian, and the compactness of $\mathcal{U}$ and $\mathcal{D}$ and continuity of $f$ guarantee that $H$ is continuous w.r.t $(x,p)$.
@@ III-B R-CLVF
In this part, we focus on the infinite-time horizon and introduce the R-CLVF.
@@
Notice the domain of the R-CLVF is $\mathcal{D}_{\gamma}$, where $\mathcal{D}_{\gamma}$ can be $\mathbb{R}^{n}$ if the limit exists on $\mathbb{R}^{n}$, while for the TV-R-CLVF, it is always $\mathbb{R}^{n}$. Also notice that for the R-CLVF, $\gamma$ is strictly greater than 0. The reason is that if $\gamma=0$, the limit in (14) may exists on a closed set.
@@
Note that though in Definition 4 from [22], we claimed the domain is a compact set; unfortunately, this is incorrect. As a result, the proof in Remark 1 in [22] is also incorrect, and we can show that the convergence is not uniform. To see these, consider a 1D toy example $\dot{x}=x+u$, where $u\in[-1,1]$ and $\gamma>0$. Then, for any trajectory starting from $x\in(-1,1)$, there exists a control signal so that the trajectory reach 00 at some finite time, therefore the limit in (14) exists. For trajectories starting from $x=\pm 1$, with best control efforts, they can only stay at the same point, so the limit in (14) does not exist. For trajectories starting from $x>1$ or $x<-1$, the trajectory will diverge, and the limit in (14) does not exist. By taking feedback
@@
$V_{\gamma}^{\infty}(x)$ is attained at $t=\min\{0,-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}\}$, and as $x\rightarrow\pm 1$, $t\rightarrow-\infty$, so the convergence is not uniform. Further, we could compute the maximum value is actually given by $|x|$ if $t=0$ and $\frac{\gamma^{\gamma}}{(1-|x|)^{\gamma}(\gamma+1)^{\gamma}}(1-\frac{\gamma}{1+\gamma})$ if $t=-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}$. Therefore, as $x\rightarrow\pm 1$, $V_{\gamma}^{\infty}(x)\rightarrow\infty$. On the other hand, when $\gamma=0$, for trajectories starting from $[-1,1]$, they could at least stay at their initial state, and the value is $V_{\gamma}^{\infty}(x)=|x|$, which is attained at the initial time. Therefore, the convergence is uniform, and the domain is closed. To avoid this situation, we reiterate that for R-CLVF, $\gamma>0$.
@@
The R-CLVF value of a state $x$ captures the maximum exponentially amplified distance between the optimal trajectory (starting from this state) and the zero-level set of $\ell(x)$. If this value is finite, the optimal trajectory converges to the SRCIS under an exponential rate no slower than $\gamma$. We will build up this conclusion via the following Theorems and Lemmas.
@@
The R-CLVF is locally Lipschitz continuous for all $x\in\mathcal{D}_{\gamma}$, and the Lipschitz constant depends on the state.
@@
See the Appendix of [33]. ∎
Further, it is not hard to see that if the domain $\mathcal{D}_{\gamma}$ is $\mathbb{R}^{n}$, then R-CLVF is radially unbounded, i.e. $\lim_{\|x\|\rightarrow\infty}V_{\gamma}^{\infty}(x)=\infty$. This is because the R-CLVF is lower bounded by a radially unbounded function $\ell$. When the domain is an open subset of $\mathbb{R}^{n}$, we have $\lim_{x\rightarrow\partial\mathcal{D}_{\gamma}}V_{\gamma}^{\infty}(x)=\infty$.
@@
The proof of this statement requires the R-CLVF-DPP (which has not been presented yet) and is provided in the Appendix of [33]. ∎
We now show that the R-CLVF satisfies the DPP, and is the viscosity solution to a variational inequality. These two Theorems are crucial for the proof of later results. They also serve as the theoretical foundation for computing the R-CLVF numerically. The proofs of these two Theorems are considered the main contributions of the paper and provide a means for constructing the R-CLVF for general nonlinear systems with bounded control and disturbance.
@@
$V_{\gamma}^{\infty}(x)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V^{\infty}_{\gamma}(z),\hskip 14.22636pt$
@@
[…]
@@
[…]
@@
(R-CLVF-VI viscosity solution) The R-CLVF is the solution to the following R-CLVF-VI in the viscosity sense,
@@
W.O.L.G, we could always assume $x$ is a strict local maxima (minima), and the maximum (minimum) value is 00, i.e., $V_{\gamma}^{\infty}(x)-\phi(x)=0$ ($V_{\gamma}^{\infty}(x)-\psi(x)=0$). As mentioned before, the definitions of the viscosity solution for TVP and IVP are different. Here, since the R-CLVF is the limit function of TV-R-CLVF, we use the definition for the TVP.
@@
Muliplly both side with $e^{\gamma(s-t)}$ and integrate on both side for $s\in[t,t+\delta]$, we get:
$e^{\gamma\delta}\phi(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))-\phi(x)\leq-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Since $V_{\gamma}^{\infty}-\phi$ has local maxima at $x$ and the maximum value is 00, we have
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
@@
Combine (25) and (26), and because $d$ is arbitrary, we have
@@
$\hskip 60.00009pte^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))\bigr{\}}$
@@
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x)\cdot f(x,u,d)+\gamma V_{\gamma}^{\infty}(x)\geq\theta_{2}>0.$ (28)
@@
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
$\geq$ $\phi(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}^{\infty}(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Again plug in DPP (3), we have
$V_{\gamma}^{\infty}(x)\geq$ $\sup_{\lambda}\inf_{u(\cdot)}e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
@@
$U(x)=\begin{cases}e^{\frac{\gamma(|x|-b)}{2}}&|x|>a\\ |x|&|x|\leq a\end{cases}$ (30)
is also one viscosity solution for (19), with any $a\gamma=2$, $b=\frac{2-2\ln{a}}{\gamma}$. The R-CLVF (29) is the viscosity solution for any $\gamma$.
Note that in Theorem 4, we do not specify the boundary condition. Therefore, the R-CLVF-VI cannot be directly solved. However, this is not a problem both theoretically and numerically. Theoretically, we only need to show that the R-CLVF satisfies this R-CLVF-VI in the viscosity sense. Numerically, we build up the solver based on the Level-set Toolbox [35], which is only applicable to time-varying PDEs (like equation (6) and (13)). Therefore, equation (13) is solved and backpropagated using DP until convergence, to get the R-CLVF [30].
Remark 5
The non-uniqueness of R-CLVF-VI does not affect the main result of this article. As will be shown later, the viscosity solution of R-CLVF-VI indicates the exponential stabilizability. Though the R-CLVF is not the unique one, we have proved that it is indeed one of the solutions.
@@
The proof can be analogously obtained from [22] and is omitted. This Proposition is vital for the proof of Lemma 6, as it states the Lie derivative of the R-CLVF along the system dynamics is less than or equal to $\gamma V_{\gamma}^{\infty}$.
III-C Existence of R-CLVF
We are now ready to present the main result of the article by introducing Lemmas 5 and 6 that will together form Theorem 7. We first show that one sufficient condition of the existence of the R-CLVF on $\mathcal{D}_{\gamma}$ is the exponential stabilizability from the ROES to the SRCIS. Later, we will show that this is also a necessary condition.

### v3

@@ III-A TV-R-CLVF
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda_{t}}\inf_{u(\cdot)\in\mathbb{U}_{t}}J_{\gamma}(t,x,u(\cdot),\lambda),$ (9)
@@
$J_{\gamma}(t,x,u(\cdot),d(\cdot))=\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\bigl(\xi(s;t,x,u(\cdot),d(\cdot))\bigl),$ (10)
@@
The TV-R-CLVF is Lipschitz in $x,t$ and bounded for any open set $\mathcal{C}$ .
@@
See the Appendix. A-A. ∎
We now show that the TV-R-CLVF satisfies the DPP and is the unique viscosity solution to the TV-R-CLVF-VI. Theorem 1 forms the foundation for the proof of Theorem 5; therefore, its proof is included in the main text of the paper.
@@
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda_{t}}\inf_{u\in\mathbb{U}}\max\biggl\{e^{\gamma\delta}V_{\gamma}(\xi(t+\delta),t+\delta),$
@@
Proof:
Let $W(x,t)$ denotes the RHS of (1). By definition of $W$, $\forall\lambda\in\Lambda_{t}$, $\exists\bar{u}(\cdot)\in\mathbb{U}_{t}$, s.t. for all $\varepsilon_{1}>0$,
$W(x,t)\geq\max\bigr\{e^{\gamma\delta}V_{\gamma}(\xi(t+\delta;t,x,\bar{u},\lambda[\bar{u}]),t+\delta),$
$\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\bar{u},\lambda[\bar{u}]))\bigl\}-\varepsilon_{1}.$ (12)
Further, by equation (9), for all $\varepsilon>0$, $\exists\bar{\lambda}\in\Lambda_{t}$ s.t. for all $u(\cdot)\in\mathbb{U}_{t}$,
$V_{\gamma}(x,t)\leq\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u,\bar{\lambda}[u]))+\varepsilon.$ (13)
Denote $y=\xi(t+\delta;t,x,\bar{u},\lambda[\bar{u}]),t+\delta)$. Again, by equation (9), $\forall\lambda\in\Lambda_{t+\delta}$, $\exists\tilde{u}\in\mathbb{U}_{t+\delta}$, s.t. $\forall\varepsilon_{2}>0$,
$V_{\gamma}(y,t+\delta)\geq\max_{s\in[t+\delta,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\tilde{u},\lambda[\tilde{u}]))-\varepsilon_{2}.$
Multiplying $e^{\gamma\delta}$ on both side, we get
$)\displaystyle e^{\gamma\delta}V_{\gamma}(y,t+\delta)\geq\max_{s\in[t+\delta,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\tilde{u},\lambda[\tilde{u}$]))
$-e^{\gamma\delta}\varepsilon_{2}.$ (14)
Define the control:
$\hat{u}(s):=\begin{cases}\bar{u}(s)\ &\text{if}\quad t\leq s<t+\delta,\\ \tilde{u}(s)\ &\text{if}\quad t+\delta\leq s\leq 0,\end{cases}$
and disturbance $\hat{\lambda}$ s.t. $\hat{\lambda}[\tilde{u}](x)=\bar{\lambda}[\tilde{u}](s)$ for all $s\in[t,t+\delta]$. Plugging (1) into equation (1):
$W(x,t)\geq$ $\max\bigr\{\max_{s\in[t+\delta,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\lambda[\hat{u}]))-e^{\gamma\delta}\varepsilon_{2},$
$\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\hat{\lambda}[\hat{u}]))\bigl\}-\varepsilon_{1},$
$\geq$ $),\displaystyle\max\bigr\{\max_{s\in[t+\delta,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\hat{\lambda}[\hat{u}$])),
$\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\hat{\lambda}[\hat{u}]))\bigl\}-\varepsilon_{1}-e^{\gamma\delta}\varepsilon_{2},$
$=$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\hat{\lambda}[\hat{u}]))\bigl\}-\varepsilon_{1}-e^{\gamma\delta}\varepsilon_{2}.$
Notice we have replaced $\bar{u},\tilde{u}$ with $\bar{u}$ and $\bar{\lambda}$ with $\hat{\lambda}$ in the corresponding time intervals. Notice also that (13) still holds with $\hat{u}$ and $\hat{\lambda}$, therefore, we have
$W(x,t)\geq V_{\gamma}(x,t)-\varepsilon_{1}-e^{\gamma\delta}\varepsilon_{2}-\varepsilon.$ (15)
On the other hand, by definition of $W$, $\forall\varepsilon_{4}>0$, $\exists\bar{\lambda}\in\Lambda_{t}$ s.t. $\forall u(\cdot)\in\mathbb{U}_{t}$,
$W(x,t)\leq\max\bigr\{e^{\gamma\delta}V_{\gamma}(\xi(t+\delta;t,x,u,\bar{\lambda}[u]),t+\delta),$
$\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u,\bar{\lambda}[u]))\bigl\}+\varepsilon_{4}.$ (16)
Similarily, denote $y=\xi(t+\delta;t,x,u,\bar{\lambda}[u]),t+\delta)$. By (9), $\forall\varepsilon_{5}>0$, $\exists\tilde{\lambda}\in\Lambda_{t+\delta}$, s.t. $\forall u(\cdot)\in\mathbb{U}_{t+\delta}$,
$V_{\gamma}(y,t+\delta)\leq\max_{s\in[t+\delta,0]}e^{\gamma(s-t-\delta)}\ell(\xi(s;t+\delta,y,u,\tilde{\lambda}[u]))+\varepsilon_{5}.$ (17)
Define the disturbance:
$\hat{\lambda}(s):=\begin{cases}\bar{\lambda}(s)\ &\text{if}\quad t\leq s<t+\delta,\\ \tilde{\lambda}(s)\ &\text{if}\quad t+\delta\leq s\leq 0.\end{cases}$
Plugging (17) into (1), we get
$W(x,t)\leq\max\bigr\{\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u,\bar{\lambda}[u])),$
$\max_{s\in[t+\delta,0]}e^{\gamma(s-t)}\ell(\xi(s;t+\delta,y,u,\tilde{\lambda}[u]))+e^{\gamma\delta}\varepsilon_{5}\bigl\}+\varepsilon_{4},$
which implies
$W(x,t)\leq\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u,\hat{\lambda}[u]))+e^{\gamma\delta}\varepsilon_{5}+\varepsilon_{4},$ (18)
Again, by (9), $\forall\lambda\in\Lambda_{t}$, $\exists\hat{u}\in\mathbb{U}_{t}$ s.t.
$).\displaystyle V_{\gamma}(x,t)\geq\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,\hat{u},\lambda[\hat{u}$])). (19)
Combining (18) and (19), we have
$W(x,t)\leq V_{\gamma}(x,t)+e^{\gamma\delta}\varepsilon_{5}+\varepsilon_{4}.$ (20)
From (15) and (20), the proof is done. ∎
@@
Proof:
See the Appendix A-B. ∎
The proof of Theorem 2 is analogous to that of Theorem 6. Since the main focus of this paper is Theorem 6, we include the proof of Theorem 2 in the Appendix. It should be noted that the definitions of viscosity solutions for the terminal value problem (TVP) and initial value problem (IVP) are slightly different. For more details, the reader is referred to [19, 18] for IPV and [9] for TVP.
With Theorems 1 and 2, the TV-R-CLVF can be obtained by solving equation (21) for $s\in[t,0]$. Note that in equation (21), define $H:\mathbb{R}^{n}\times\mathbb{R}^{n}\mapsto\mathbb{R}$,
$H(x,p)=\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}p\cdot f(x,u,d).$ (22)
$H$ is called the Hamiltonian, and the compactness of $\mathcal{U}$ and $\mathcal{D}$ and Lipschitz continuity of $f$ guarantee that $H$ is bounded and Lipschitz continuous with respect to $(x,p)$.
@@ III-B R-CLVF
In this section, we introduce the R-CLVF. We will show one sufficient condition for the existence of the R-CLVF, and show some key properties of it. Finally, we will also show that the R-CLVF is a viscosity solution to a VI.
@@
Notice the domain of the R-CLVF is $\mathcal{D}_{\gamma}$, where $\mathcal{D}_{\gamma}$ can be $\mathbb{R}^{n}$ if the limit exists on $\mathbb{R}^{n}$, in contrast, for the TV-R-CLVF, the domain is always $\mathbb{R}^{n}$. Also notice that for the R-CLVF, $\gamma$ is strictly greater than 0. The reason is that if $\gamma=0$, the limit in (23) may exist on a closed set.
@@
Note that though in Definition 4 from [29], we claimed the domain is a compact set, this is incorrect. As a result, the proof in Remark 1 in [29] is also incorrect, and we can show that the convergence is not uniform. To see these, consider a 1D toy example $\dot{x}=x+u$, where $u\in[-1,1]$ and $\gamma>0$. Then, for any trajectory starting from $x\in(-1,1)$, there exists a control signal so that the trajectory reaches 0 at some finite time, therefore the limit in (23) exists. For trajectories starting from $x=\pm 1$, with best control efforts, they can only stay at the same point, so the limit in (23) does not exist. For trajectories starting from $x>1$ or x<1x<-1, the trajectory will diverge, and the limit in (23) does not exist. By taking feedback
@@
$V_{\gamma}^{\infty}(x)$ is attained at $t=\min\{0,-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}\}$, and as $x\rightarrow\pm 1$, $t\rightarrow$-$\infty$, so the convergence is not uniform. Further, we could compute the maximum value is actually given by $|x|$ if $t=0$ and $\frac{\gamma^{\gamma}}{(1-|x|)^{\gamma}(\gamma+1)^{\gamma}}(1-\frac{\gamma}{1+\gamma})$ if $t=-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}$. Therefore, as $x\rightarrow\pm 1$, $V_{\gamma}^{\infty}(x)\rightarrow\infty$. On the other hand, when $\gamma=0$, for trajectories starting from [1,1][-1,1], they could at least stay at their initial state, and the value is $V_{\gamma}^{\infty}(x)=|x|$, which is attained at the initial time. Therefore, the convergence is uniform, and the domain is closed. To avoid this situation, we reiterate that for R-CLVF, $\gamma>0$.
For different choices of $\gamma$, do we obtain completely different R-CLVFs? Although the function values can vary significantly with $\gamma$, we can show that for any $\gamma>0$, the zero sub-level set remains unchanged. Denote the zero sub-level sets of TV-R-CLVF and R-CLVF as
$\mathcal{Z}_{\gamma}(t):=\{x:V_{\gamma}(x,t)\leq 0\},\hskip 10.00002pt\mathcal{Z}_{\gamma}^{\infty}:=\{x:V_{\gamma}^{\infty}(x)\leq 0\}.$
Lemma 3
Proof:
$\ell(\xi(t_{1};t,x,u_{1}(\cdot),\lambda[u_{1}])<0$
for some $t_{1}$. On the other hand, since $V_{\gamma_{2}}(x,t)>0$ and $e^{\gamma_{2}(s-t)}>0$, there exists $\bar{\lambda}$ and for all $u(\cdot)$ s.t.
$\ell(\xi(t_{1};t,x,u(\cdot),\bar{\lambda}[u])\geq 0$
for all $s\in[t,0]$. However, applying $u_{1}(\cdot)$ gives
$\ell(\xi(t_{2};t,x,u_{1}(\cdot),\bar{\lambda}[u_{1}])<0$
for some $t_{2}$, which is a contradiction. Therefore, $V_{\gamma_{1}}(x,t)<0$ implies $V_{\gamma_{2}}(x,t)\leq 0$. Switching $\gamma_{1}$ and $\gamma_{2}$, we get the same result, and therefore we have $V_{\gamma_{1}}(x,t)\leq 0$ if and only if $V_{\gamma_{1}}(x,t)\leq 0$. With the same process, we could also show that $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$ if and only if $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$. Further, the above inequalities also hold when either $\gamma_{1}$ or $\gamma_{2}$ is 0, and when $\gamma=0$, in the infinite-time horizon, the zero sub-level set is the SRCIS. Therefore, we have proved $\mathcal{Z}_{\gamma}^{\infty}=\mathcal{I}_{m}$.
This lemma links the R-CLVF to the SRCIS by showing that the SRCIS is exactly the zero sublevel set of the R-CLVF. Since our goal is to prove that an R-CLVF implies exponential stabilizability to the SRCIS with rate $\gamma$, it is essential that the R-CLVF encode the SRCIS in a $\gamma$-independent way. In particular, although the SRCIS itself does not depend on $\gamma$, the R-CLVF values do. We therefore establish that for any $\gamma>0$, the R-CLVF shares the same zero sublevel set, and that set is exactly the SRCIS.
Since the R-CLVF is defined as the limit function of TV-R-CLVF, it is natural to ask under what conditions the limit exists. Consider an initial state $x$ in the SRCIS, then $\ell(x)<0$. Then, for its value to be 0, $\ell(\xi(s;t,x,u(\cdot),d(\cdot)))$ has to converge to 0 with an exponential rate $\gamma$, which cannot be guaranteed. Therefore, we have $V_{\gamma}^{\infty}(x)\leq 0$ for all $x\in\mathcal{I}_{m}$, regardless of $\gamma$. This means the limit always exists in the SRCIS.
For the states outside $\mathcal{I}_{m}$, the value can be infinity because of the exponential amplifier $\gamma$. The next Lemma shows that the exponential stabilizability to SRCIS is a sufficient condition for the existence of the R-CLVF.
Lemma 4 (Existence of R-CLVF)
The R-CLVF exists on an open set $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$) if the system is exponentially stabilizable (under rate $\gamma$) to its SRCIS from $\mathcal{D}_{\text{ROES}}$ (or $\mathbb{R}^{n}$), despite worst-case disturbance. Further $\mathcal{D}_{\text{ROES}}\subset\mathcal{D}_{\gamma}$.
Proof:
Assume the system is exponentially stabilizable to the SRCIS. Using Definition 2, we have $\forall\lambda\in\Lambda$, $\exists u^{*}(\cdot)\in\mathbb{U}$, $\exists c>0$ s.t.
Plugging in equation (2),
Plugging in $\ell(x)=h(x)-V^{\infty}_{m}=\|x\|-V^{\infty}_{m}$, we have
$\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))$
$=$ $h(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))-V^{\infty}_{m},$
$=$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\max_{a\in\partial\mathcal{I}_{m}}\|a\|,$
and
$\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\min_{a\in\partial\mathcal{I}_{m}}\|a\|,$
$=$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl(\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\|a\|\bigl),$
$\leq$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl(\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|\bigl),$
$\leq$ $ce^{-\gamma(s-t)}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|,$ (25)
where we used equation (4) for the last inequality. Multiplying $e^{\gamma(s-t)}$ on both side
which holds for all $s\in[t,0]$. Therefore
$V_{\gamma}(x,t)=$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\left(\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))\right),$
$\leq$ $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$
This upper bound $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$ is independent of $t$, therefore as $t\rightarrow$-$\infty$, we have $V_{\gamma}^{\infty}(x)\leq c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$. Since the R-CLVF monotonically increases, we conclude that the limit in (23) exists $\forall x\in\mathcal{D}_{\text{ROES}}$. This also implies $\mathcal{D}_{\text{ROES}}\subset\mathcal{D}_{\gamma}$.
@@
This Lemma shows that for nonlinear systems, the existence of the R-CLVF can be justified by analyzing the system dynamics. This sufficient condition is intuitive. The R-CLVF value of a state $x$ captures the maximum exponentially amplified distance between the optimal trajectory (starting from this state) and the zero-level set of $\ell(x)$. For this value to be finite, the optimal trajectory must converge to the SRCIS under an exponential rate no slower than $\gamma$. In Section III-C, we will show that this is also a necessary condition.
@@
The R-CLVF is locally Lipschitz continuous for all $x\in\mathcal{D}_{\gamma}$, if $\gamma\leq\alpha$, where $\alpha$ is the fastest rate of exponential stabilizability of the system.
@@
See the Appendix A-C . ∎
Further, it can be seen that if the domain $\mathcal{D}_{\gamma}$ is $\mathbb{R}^{n}$, then R-CLVF is radially unbounded, i.e. $\lim_{\|x\|\rightarrow\infty}V_{\gamma}^{\infty}(x)=\infty$. This is because the R-CLVF is lower bounded by a radially unbounded function $\ell$. When the domain is an open subset of $\mathbb{R}^{n}$, we have $\lim_{x\rightarrow\partial\mathcal{D}_{\gamma}}V_{\gamma}^{\infty}(x)=\infty$.
@@
See the Appendix A-D. ∎
We now show that the R-CLVF satisfies the DPP, and is the viscosity solution to a VI. These two Theorems are crucial for the proof of later results. The proofs of these two Theorems are considered the main contributions of the paper and provide a means for constructing the R-CLVF for general nonlinear systems with compact control and disturbance.
@@
$V_{\gamma}^{\infty}(x)=\sup_{\lambda\in\Lambda_{t}}\inf_{u\in\mathbb{U}}\max\biggl\{e^{\gamma\delta}V^{\infty}_{\gamma}(z),\hskip 14.22636pt$
@@
$V_{\gamma}^{\infty}(x)\leq V_{\gamma}(x,t)+\varepsilon=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl\{$
@@
$V_{\gamma}^{\infty}(x)\geq V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl\{$
@@
(R-CLVF-VI) The R-CLVF is a solution to the following R-CLVF-VI in the viscosity sense,
@@
Without loss of generality, we could always assume $x$ is a strict local maxima (minima), and the maximum (minimum) value is 0, i.e., $V_{\gamma}^{\infty}(x)-\phi(x)=0$ ($V_{\gamma}^{\infty}(x)-\psi(x)=0$). Here, since the R-CLVF is the limit function of TV-R-CLVF, we use the definition for the TVP.
@@
Multiplying $e^{\gamma(s-t)}$ on both side and integrating on both side for $s\in[t,t+\delta]$, we get:
$e^{\gamma\delta}\phi(\xi(t+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))-\phi(x)\leq-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Since $V_{\gamma}^{\infty}-\phi$ has local maxima at $x$ and the maximum value is 0, we have
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(t+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
@@
Combining (36) and (37), and because $d$ is arbitrary, we have
@@
$\hskip 60.00009pte^{\gamma\delta}V_{\gamma}^{\infty}(\xi(t+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))\bigr\}$
@@
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\psi(x)\cdot f(x,u,d)+\gamma V_{\gamma}^{\infty}(x)\geq\theta_{2}>0.$ (39)
@@
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(t+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
$\geq$ $\psi(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}^{\infty}(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Again plugging in DPP (5), we have
$V_{\gamma}^{\infty}(x)\geq$ $\sup_{\lambda}\inf_{u(\cdot)}e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(t+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}])),$
@@
We do not have the uniqueness guarantee for two reasons: 1) we do not specify any boundary conditions, and 2) the spatial growth of the R-CLVF is not bounded, as shown in Proposition 3. Further, as will be seen in the next paragraph, it is not trivial to specify the boundary condition.
@@
$U(x)=\begin{cases}e^{\frac{\gamma(|x|-b)}{2}}&|x|>a\\ |x|&|x|\leq a\end{cases}.$ (41)
is also one viscosity solution for (30), with any $a\gamma=2$, $b=\frac{2-2\ln{a}}{\gamma}$. Also, $U$ also satisfies that $U(x)=0$ at SRCIS, and $U(x)\rightarrow\infty$ as $x\rightarrow\infty$. This means specifying boundary conditions on the SRCIS or on the boundary of $\mathcal{D}_{\gamma}$ cannot enforce the uniqueness of the viscosity solution. The R-CLVF (40) is the viscosity solution for any $\gamma$.
This means that if we directly solve the R-CLVF-VI, we might not get the R-CLVF as desired. However, this is not a problem both in theory and for the numerical computation. In theory, implies Proposition 4, which is vital to prove one main result (Theorem 8). For numerical computation, we build up the solver based on the Level-set Toolbox [36], which is only applicable to time-varying PDEs (like equation (6) and (21)). More specifically, we discretize the state space, solve (21) and backpropagate using DP until convergence to obtain (23), instead of directly solving (30).
@@
Proof:
See the Appendix A-E
This Proposition is vital for the proof of Lemma 7. Also note that in the proof, the super-differential and sub-differential are used for non-differentiable points.
III-C R-CLVF implies Exponential Stabilizability
As we have already shown that one sufficient condition of the existence of the R-CLVF is the exponential stabilizability in Lemma 4, we will now show that this is also a necessary condition.